The Knocking at the Door :

A Fantasy on Fate, Forster and Beethoven’s Fifth

BY

Anne FOATA

Near the beginning of Howards End,^{1} the novel’s two heroines, Margaret and Helen Schlegel attend a concert at the Queen’s Hall in London. They don’t think much of the Brahms, Mendelssohn and Elgar pieces on the programme, but Beethoven’s Fifth Symphony has their full attention, while their younger brother Tibby follows its performance from the “full score open on his knee” (p. 43). Helen has a knack for finding meanings in music, for “turning” it into literature or painting. For Beethoven’s Symphony she devises this startling fantasy, which has since become Forster’s much anthologized tour de force.

“[T]he Andante had begun — very beautiful, but bearing a family likeness to all the other beautiful Andantes that Beethoven has written, and, to Helen’s mind, rather disconnecting the heroes and shipwrecks of the first movement from the heroes and goblins of the third. She heard the tune through once, and then her attention wandered […] and the Andante came to an end… Helen said to her aunt : ‘Now comes the wonderful movement : first of all the goblins, and then a trio of elephants dancing,’ and Tibby implored the company generally to look out for the the transitional passage on the drum…

‘No ; look out for the part where you think you have done with the goblins and they come back,’ breathed Helen, as the music started with a goblin walking quietly over the universe, from end to end. Others followed him. They were not aggressive creatures ; it was that that made them so terrible to Helen. They merely observed in passing that there was no such thing as splendour or heroism in the world. After the interlude of elephants dancing, they returned and made the observation for the second time. Helen could not contradict them, for, once at all events, she had felt the same, and had seen the reliable walls of youth collapse. Panic and emptiness ! Panic and emptiness ! The goblins were right.

ANNE FOATA

Her brother raised his finger : it was the transitional passage on the drum.

For, as if things were going too far, Beethoven took hold of the goblins and made them do what he wanted. He appeared in person. He gave them a little push, and they began to walk in a major key instead of in a minor, and then he blew with his mouth and they were scattered ! Gusts of splendour, gods and demigods contending with vast swords and fragrance broadcast on the field of battle, magnificent death ! […] Any fate was titanic ; any contest desirable ; conqueror and conquered would alike be applauded by the angels of the utmost stars.

And the goblins – they had not really been there at all ? They were only the phantoms of cowardice and unbelief ? One healthy human impulse would dispel them. Men like the Wilcoxes, or President Roosevelt, would say yes. Beethoven knew better. The goblins really had been there. They might return – and they did. It was as if the splendour of life might boil over and waste to steam and froth. In its dissolution one heard the terrible, ominous note, and a goblin, with increased malignity, walked quietly over the universe from end to end. Panic and emptiness ! Panic and emptiness ! Even the flaming ramparts of the world might fall.

Beethoven chose to make all right in the end. He built the ramparts up. He blew with his mouth for the second time, and again the goblins were scattered. He brought back the gusts of splendour, the heroism, the youth, the magnificence of life and death, and, amid vast roarings of a superhuman joy, he led his Fifth Symphony to its conclusion. But the goblins were there. They could return. He had said so bravely, and that is why one can trust Beethoven when he says other things.

Helen pushed her way out during the applause… The music had summed up to her all that had happened or could happen in her career. She read it as a tangible statement, which could never be superseded. The notes meant this and that to her, and they could have no other meaning, and life could have no other meaning (chap. V ; p. 45–47 ; my emphasis).”

Helen’s interpretation of Beethoven’s Fifth Symphony not only, as she thinks, sums up her own life and her vision of life in general ; it is my contention that it points to the very structure of Howards End, by highlighting the four significant moments or episodes that compose it. It thus provides the musical subtext to the novel, its necessary complement.

FATE, FORSTER AND BEETHOVEN’S FIFTH

To paraphrase Forster’s own beginning, I may as well begin with the epigraph of the novel, which reads “Only connect…,” and state summarily that Forster’s cautionary words to his characters, his readers, and the world at large, contain in a nutshell the substance of the novel. What there is to “connect,” are, notably, the two antithetical worlds of business and culture, embodied by the Wilcox family (or to be more precise, the male Wilcoxes) and the Schlegels. The Wilcoxes “have their hands on all the ropes” ; they are the great English industrialists of the first decade of the XXth century, the Empire builders, efficient, ruthless, indifferent to the masses they employ and exploit. The leisured intellectual Schlegel sisters try to educate those “obtuse” Wilcoxes, make them acknowledge the importance of “personal relations” in life. They devote themselves to finding some sort of working “connexion” with them, at the same time striving to establish yet another “connexion” with the world of the Basts. Leonard Bast and his wife Jacky represent the “submerged,” “those who are obliged to pretend that they are gentlefolk” (p. 58), though they are poor enough to actually harbour a suspicion that ladies like the Schlegel sisters might steal their neighbours’ umbrellas during concerts. The Basts thus provide the field of experimentation for both the Schlegel and Wilcox worlds and eventually the stumbling block that proves them both wanting. Some malignant power, indeed, seems to lie forever in wait to undo the efforts of men and women of good will ; let’s call it Fate.

Here we come back to Beethoven’s Fifth Symphony and to Helen’s fanciful interpretation of it. “So pocht das Schicksal an die Pforte,” Beethoven had told his friend and secretary Schindler to explain the famous first four notes of the Overture. The initial theme of Fate (“knocking at the door”) is to remain the dominant note throughout the symphony, running now subdued, now triumphant, through its four movements, man and Fate locked in a “titanic” contest until the former’s blazing victory in the Finale. Helen’s vision of the Finale is less reassuring, though ; some victory has been achieved, it remains with each reader to assess its quality.

Helen’s personification of Fate is a goblin, an invading army of goblins^{2} “walking quietly over the universe, from end to end.” They are not aggressive, “[t]hey merely observe in passing that there [is] no such thing as splendour or heroism in the world.” They are indifferent to men ; it is that that makes them so terrible.

Helen’s grappling with the goblins, her attempts to crack their damaging indifference, in other words, to make them “disconnect” from mankind and allow it its share of splendour, has to do with her dealings with both Basts and Wilcoxes. Fate for her bears a name and a face : Henry Wilcox’s, the man who wronged Leonard in the past by seducing Jacky and turning her into a kept woman, the man who will give bad professional advice and eventually be responsible for the couple’s destitution. It is worth noting that a reference to a goblin or a “goblin footfall,” meaning Fate’s blind treading over one’s life, occurs at three different points in the narrative, which are also crucial moments for Leonard’s desperate attempts at surfacing from his squalid depths (p. 57 ; 122 ; 315).

Helen’s fighting with the Wilcoxes, on her behalf and over the Basts, has the somewhat unbalanced edge of her own high-strung romantic personality. Margaret is more mature. Her crusade against Fate and the crass Wilcoxes, that is, her own attempts at a connexion with them and at challenging them to an awareness of their responsibilities, shows more “proportion,” that Greek ideal passed on to Forster by one of his “lawgivers,” Montaigne.^{3} In fact Margaret appears as Montaigne’s quintessentially “honnête homme,”^{4} the individual endowed with reason and good will. In Helen’s analysis of the Fifth Symphony she may actually be Beethoven himself who “[blows] with his mouth” and scatters the goblins, or “[takes] hold of [them] and [makes] them do what he [wants].

Beethoven’s first movement is described by Helen as having to do with “heroes and shipwrecks.” This rather terse statement, however, in view of her sustained lyrical ravings in the third, is more than adequate information on what constitutes for her a sharp moment of awareness and the novel’s prologue (chap. I–IV). The Allegro con brio sets the pace of her enthusiastic discovery of the Wilcoxes during a brief visit at Howards End, their country house in Hertfordshire. Henry Wilcox bowls over all her cherished ideals, and one by one all the “Schlegel fetishes,” Art and Literature, Women’s Equal Rights, Liberalism and Tolerance, Civility towards one’s menials and Responsibility to the poor, get overthrown like as many ninepins. Helen is “stunned,” but “fascinated” by the Wilcoxes. Theirs must be the real world, the great outer world of heroic enterprise and efficiency. To share their life, to go motoring with them through the countryside appears to her as “the supreme joy of life.” When the younger son Paul comes home from school, “ready to flirt with any pretty girl,” she [meets] him halfway” within the “halo of Romance” she has built around the whole family ; they kiss in the moonlight under a magic wych-elm. Next morning at breakfast, surrounded by the family, he looks frightened and Helen’s romantic construction collapses. The Wilcoxes after all are “a fraud,” “just a wall of newspapers and motor-cars and golf-clubs” with nothing behind but “panic and emptiness” (p. 40).

This short episode of a single week in June is to set Helen’s erratic behaviour throughout the following four years or so covered by the novel and to steer the plot through the turbulence of its main part (the Scherzo part) to the reversal of situation in its finale. Indeed, the ever present narrator warns the reader that “[t]o Helen, at all events, her life was to bring nothing more intense than the embrace of this boy (Paul) who played no part in it.” (p. 39). The kiss and the following disillusion unsettle her ; they pave the way to her high-strung, somewhat hysterical attitude towards Leonard Bast and her near-neurotic loathing of the Wilcoxes. There is no splendour or heroism in “the great outer life,” only “panic and emptiness.” The cultured Schlegels’ first attempt, through Helen,” at a “connexion” with the world of business has foundered. “Heroes and shipwrecks” indeed !

The second great moment of Howards End belongs to Margaret (chap. IV–XIII). Helen is physically absent, visiting her German cousins in Stettin, just as she is mentally absent during the concert, her attention wandering to the audience, the architecture and the “vapid” Cupids that adorn the ceiling of Queen’s Hall. It must be noted, however, that before she escapes to Germany to put some distance between herself and her first shipwreck, she has had time to initiate a second one, in her failure through flippancy and lack of self-control to secure a meaningful connexion with her neighbour at the concert, Leonard Bast. The missed opportunity reverberates through the following action “as a goblin footfall, as a hint that all is not for the best in the best of all possible worlds, and that beneath these superstructures of wealth and art there wanders an ill-fed boy who has recovered his umbrella indeed, but who has left no address behind him, and no name” (end of chap. V).

The stately movement of the Andante well suits the quiet dignity and hopeful tenor of this second climactic moment of the novel : the development, during the months of November and December of the same year, of the deep understanding between the two impressive women in Howards End. Neither Ruth Wilcox, Henry’s wife, nor Margaret can be described, at least outwardly, as strong personalities ; neither is aggressive, overbearing, overassuming. Their essential kinship does not even derive from their succeeding each other in matrimony with Henry ; it bears upon their being the two women of good will in the novel, capable of thwarting the (mis)dealings of Fate. Each of them, in a different capacity, is the recipient of what in Forster’s view are the enduring values of England. They are the representatives of the two social classes that may still (in 1910) save it from encroaching evils : the intellectual aristocracy and the yeomanry. As in Beethoven’s second movement, Fate is held in abeyance, hope is possible, but the episode ends on another “shipwreck,” the most final of all, Ruth Wilcox’s sudden death before she can take Margaret to the holy of holies, the place of salvation, Howards End. Her spirit, however, lives on.

Beethoven’s Scherzo, Helen’s “wonderful movement” of stalking goblins and dancing elephants, covers an extended one-year-period following a narrative ellipsis of more than two years. It takes up the bulk of the novel (chap. XII-XLIV). Helen’s fanciful “trio of elephants dancing” must be the trio fugato that opens the movement and is heard at intervals throughout, truncated and finally pianissimo on the clarinets, violins and oboes. In Helen’s interpretation of the Scherzo, the trio seems part of Fate’s challenge ; it sounds the ominous return of the goblins, denying mankind its share of splendour and romance.

In the Schlegels’ dealings with Leonard Bast, the goblins’ footfalls are heard after each of their endeavours to pull him out of the drabness of his life : first after his night of “heroic” wandering in the woods around London (chap. XIII & XIV) and their passing on to him Mr. Wilcox’s information about the threatening insolvency of his employers (which proves wrong, in chap. XVI), later after the momentous night at Oniton following Evie Wilcox’s Shropshire wedding (chap. XXVII and parts of XL & XLI). After each romantic episode, with Helen striving to build up his ego and almost succeeding in making him believe in the Great Adventure of Life (trying, Forster writes on page 237, “to cut the rope that fastened [him] to the earth”), Leonard’s fall gets heavier and his rebound shorter, with any hope of recovery definitively curtailed after the Oniton episode. Helen may well appear as a goddess to poor earth-bound Leonard, it nevers occurs to her that her flamboyant, erratic doings combine with Henry Wilcox’s bad piece of professional advice and former indiscretion with Jacky to push him further down into destitution and eventual death. In the whole of this third episode of Howards End, Helen furthermore appears as her own self-appointed Nemesis as she crushes herself together with Leonard under the weight of her high-strung Romanticism and her somewhat unbalanced way of living up to it. Both her near-neurotic refusal to “connect” with the world of the Wilcoxes and her wild endeavours on behalf of the Basts are apt to appear as sheer bungling ; as finally does her sole “successful” attempt, which results in the haphazard conception of their son. This is not one of the least cosmic jokes that Fate in accordance with the etymological meaning of the word scherzo (from the Italian scherzare : to joke) has been playing throughout the narrative of this third momentous period of Howards End. Quite consistent also with the mischievious atmosphere of both the scherzo movement and the goblins’ misdeeds is Leonard’s pitiful but ludicrous death at the hand of Henry Wilcox’s son Charles, felled as he is by a weak heart and the blunt edge of the Schlegel sisters’ father’s Prussian sword which had no reason at all to hang unsheathed on the wall at Howards End. The joke has turned into tragedy ; whatever form Fate may care to assume, stalking goblins or dancing elephants, it certainly knows how to crush its quarry.

The scattering of the goblins, when it occurs in this third movement of the novel, is Margaret’s doing. Fastidious but sensible Margaret sets about to “connect the prose and the passion” of life by accepting to marry Henry Wilcox against the wild outbursts of Helen, thus joining by the holy bonds of matrimony both the world of liberal intelligentsia and that of exploitive capitalism. To Helen, Margaret appears as a “heroine,” but to Margaret at this stage the real heroes are men like Henry Wilcox, the men who made England what it is and allow leisured and cultured people like the Schlegels to enjoy their leisure and their culture. Margaret’s spirited encomium, from the heights of the Purbeck Hills near the coastal resort of Swanage, deserves full quotation.

“If Wilcoxes hadn’t worked and died in England for thousands of years, you [Helen] and I couldn’t sit here without having our throats cut. There would be no trains, no ships to carry us literary people about in, no fields even. Just savagery. No – perhaps not even that. Without their spirit life might never have moved out of protoplasm. More and more do I refuse to draw my income and sneer at those who guarantee it” (chap. XIX ; p. 177–178).”

Margaret’s prosaic engagement to Henry bears a touch of heroism, considering the gap that separates their two worlds. We are made to hear some “gusts of splendour,” to acquiesce with Beethoven that “any fate [is] titanic” indeed, “any contest desirable.” The goblins seem scattered for a while, but in no time they are seen gathering again and sounding their ominous note. At Oniton, at his daughter’s wedding, Henry’s former affair with Jacky is accidentally exposed together with his obtuse adherence to the double standard and his dogged skirting of his responsibility. Magnanimous Margaret forgives, and marries him a month later. “And the goblins – they had not really been there at all ? Men like the Wilcoxes … would say yes. But Beethoven knew better. The goblins really had been there. They might return,” and they do, with a vengeance. The overwhelming wreckage that concludes this third part of the novel’s plot undoubtedly attests to it : Leonard Bast is lying dead at Howards End ; Helen, socially disgraced by her unwedded pregnancy, has exiled herself to Germany ; Henry Wilcox callously refuses to see any similarity between Helen’s “sin” and his own indiscretion with Jack Bast and just as crassly denies his share of responsibility in Leonard’s destitution ; Charles Wilcox will be indicted for manslaughter ; Margaret decides to leave Henry and go live with Helen in Germany. In the general havoc of the scene, a “terrible, ominous note” can indeed be heard and a sneering goblin be seen walking “with increased malignity” over the universe “from end to end.” The light playful tone which is the generic character of the scherzo seems to fail at last ; the comic chassé-croisé that Fate had set them all dancing around the bewitched Howards End has come to a halt in front of the dead body of Leonard Bast.

A moment in June, some fourteen months later : the epilogue of the novel (chap. XLIV) has all the characters gathered in the shelter of Howards End. Margaret has “picked up the pieces” and out of “the black abyss of the past” tries to build them a future. Helen has born a son in the “central room” of the house, Henry Wilcox, broken by his son’s indictment and emprisonment, has settled in the country and lets himself be nursed by a forgiving wife. As in the Symphony’s Finale, the goblin theme of the scherzo tries one last time to assert its rights, hinting at a world that is not all that it should be, and indeed “Leonard was dead, Charles had two more years in prison” (p. 326), and Paul Wilcox’s hostility towards the Schlegels has not abated. But, for Helen, and undoubtedly for Forster, Margaret is “heroic” in her own way ; her “life has been heroic.” Like Beethoven, in the Finale of his Symphony, she has “brought back the gusts of splendour, the heroism, the youth, the magnificence” of a life in the country, lived in harmony with the seasons.

The epilogue of Howards End does not, however, echo “the vast roarings of a superhuman joy” with which Beethoven brings his Symphony to its conclusion. The enthusiasm is subdued : the very springs of life that were once expected to flow so liberally have contracted and exuberant vitality has shrunk to a modicum of survival. Helen has a son, but she proclaims herself cured of love, “ended,” and unable to love again ; Henry has become an invalid, Margaret herself seems strangely aloof, suffused with the witch-like wisdom that permeates the house and was its previous (female) owners’ birthright. They have settled down to a new life and Margaret-Beethoven has built the ramparts up around them. No doubt she has triumphed over the goblins and achieved the connexion that was her initial purpose, but her victory has exacted a high price from all the characters.

The last picture in Howards End of a pacified family group serenely conversing outdoors in the country while “the big meadow” is being cut, of a life “obscure, yet gilded with tranquillity” (p. 326), and attuned to the passage of the seasons, will remind the readers of yet another Symphony : Beethoven’s Sixth, the Pastoral. Similarities of intention may not appear fortuitous since their composition was strictly contemporaneous and their first public performance rendered on the same day, December 22, 1808 during a concert that Beethoven thought was to be his farewell to Vienna. The Pastoral (man’s contest with Nature and his final victory) ends on a peaceful movement that Beethoven himself entitled “Frohe and dankbare Gefühle nach dem Sturm.” These “happy and grateful feelings” are those of the Schlegel-Wilcox family finally at peace with each other after weathering the storm.

There is peace indeed, but no jubilation at the conclusion of Howards End. I wonder if Forster, like Virgil in his Eclogues, was not in fact throwing some apotropaic morsel to the monster that was to devour the countryside and forever change the life and customs of England. Both Helen and Margaret have misgivings about the abiding power of their haven of Howards End. “London is creeping,” says Helen (p. 329), Hertfordshire, Surrey, “and even Hampshire” are turning into “Suburbia” (Forster actually uses this term that sounds so modern). Pollution by the fumes of “stinking” motor cars (another recurring theme in the novel) is spreading, and, as both sisters know, “London is only part of something else” : “[l]ife’s going to be melted down, all over the world” and the “civilization of luggage” will take over. One has to remind oneself quite so often that these rueful forecasts were made as early as 1910. That they did come true needn’t be debated here, as Forster himself was to learn much to his sorrow when he considered buying back his childhood home of Rooksnest, which had been the model for Howards End.

Like Helen who “turned” music into literature, Forster “turned” Beethoven’s Fate Symphony into a piece of narrative called Howards End. But can the responsive reader quite escape the fallacy surrounding the author-narrator relationship and not make Helen’s qualified assessment Forster’s own ? The goblins have indeed been scattered. But they can return. And, seen from the vantage point of our declining century, they certainly did. Forster-Beethoven “had said so bravely, and that is why one can trust [him] when [he] says other things.”^{5}

UNIVERSITY OF STRASBOURG II (FRANCE)

NOTES

1. London : Edward Arnold, 1910. The quotations and their page references are those of the 1989 Penguin Books edition.

2. Helen’s father, a self-exiled Prussian, may have told his daughters German folktales about nasty goblins, and Forster may have been reminded of Christina Rossetti’s cruel goblins in Goblin Market.

3. In “What I believe,” Two Cheers for Democracy (London : Edward Arnold, 1951), p. 77.

4. I won’t argue here Margaret’s touch of masculinity (or Forster’s touch of feminine sensibility). She may well have been inspired by one of Goldsworthy Lowes Dickinson’s sisters, as Forster himself acknowledged, or by one of the Stephens sisters as Wilfred Stone has it (cf. Oliver Stallybrass’s introduction to the Penguin edition). She undeniably appears as a fastidious, liberal, tolerant, well-read female alter ego of Forster himself.

5. My own references to Beethoven’s Fifth were supported by Jean and Brigitte Massin’s analysis in Ludwig van Beethoven (Paris : Club du Livre Français, 1955, p. 153–178 and 666–668.)